[figure: 9.7 cm]

VISUALIZATION OF THE ANTIFERROMAGNETIC INSULATOR - FERROMAGNETIC METAL PHASE TRANSITION IN THE MANGANITE Nd_{0.5}Sr_{0.5}MnO_{3}

S.L. Gnatchenko^{1}, A.B. Chizhik^{1}, I.O. Shklyarevskiy^{1}, D.N. Merenkov^{1},

V.I. Kamenev^{2}, Yu.G. Pashkevich^{2}, K.V. Kamenev^{3}, G. Balakrishnan^{3}, D. McK Paul^{3}

1 Institute for Low Temperature Physics & Engineering, 310164 Kharkov, Ukraine

2 Donetsk PhysTech, 310114 Donetsk, Ukraine 3 Department of Physics, University of Warwick, Coventry CV4 7AL, England

It has been found that the first order phase transition from an antiferromagnetic (AFM) metallic state to a ferromagnetic (FM) insulating state in the manganite Nd_{0.5}Sr_{0.5}MnO_{3} is accompanied by a significant change of reflectivity of visible light. The reflection coefficient of the FM phase is approximately twice as large as the coefficient of the AFM phase for light with a wavelength of 633 nm. This effect has been used for the visualization of the AFM-FM phase transition in Nd_{0.5}Sr_{0.5}MnO_{3}. The coexistence of AFM and FM phases was visually observed during the spontaneous and field-induced AFM-FM transition. In both cases, the transition occurred by nucleation and expansion of domains of the energetically favorable phase. However, the periodic domain structure of magnetic intermediate state was not observed. The H − T phase diagram of Nd_{0.5}Sr_{0.5}MnO_{3} manganite was constructed using the visual observations and reflectivity measurements in fields up to 40 kOe. It should be noted that a stripe domain structure was observed in the AFM state whereas the FM state was observed to be uniform.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : S.L. Gnatchenko

Address for correspondence : Institute for Low Temperature Physics & Engineering, 47 Lenin Ave., 310164 Kharkov, UKRAINE

Email address : gnatchenko@ilt.kharkov.ua